The final exam also had a normal distribution, but with mean of 150 and standard deviation of 15. At least what score should Wright get in order to keep the same ranking (i.e , top 10 percent).
Answer by ewatrrr(24785)

mean of 150 and standard deviation of 15
z = invNorm(.90) = (x -150)/15
15invN0rm(.90) + 150 = x
......
Using a TI calculator 0r similarly a Casio fx-115 ES plus 0r NORMSINV(.90)
15(1.282) + 150 = x = 169.23
least score for top 10% = 170